OpenStudy (anonymous):

Consider the line that passes through the points (2, 3) and (6, 11). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y -11 = 2(x - 6) y - 3 = 2(x - 2) Only equation #1 represents the line that passes through the two given points. Only equation #2 represents the line that passes through the two given points. Both equation #1 and equation #2 represent the line that passes through the two given points. Neither equation #1 nor equation #2 represents the line that passes through the two given points.

OpenStudy (anonymous):

slope of line is y2-y1/x2-x1=11-3/6-2 =8/4=2 line equation is y-3=2(x-2) or y -11 = 2(x - 6) y -11 = 2(x - 6)

OpenStudy (anonymous):

A. B. C. D. ???

OpenStudy (anonymous):

Both equation #1 and equation #2 represent the line that passes through the two given points.

OpenStudy (anonymous):

Thanks